OpenStudy (anonymous):

can someone help me with this problem.. samuel pulls two coins out of his pocket randomly without replacement. if his pocket contains one nickel, one dime, and one quarter, what is the probability that he pulled more than 20 cents out of his pocket? justify your work.

OpenStudy (mertsj):

If he pulled out more than 20 cents, he had to get the quarter. The probability that the first pick was not the quarter is 2/3 The probability that the second pick was not the quarter is 1/2 So the probability he did not get the quarter is 2/3 times 1/2 = 1/3 So the probability he did get the quarter is 1-1/3=2/3

OpenStudy (anonymous):

is the answer above unclear?

OpenStudy (anonymous):

there are only 3 possibilities quarter and dime quarter and nickel dime and nickel two out of those three amount to more than 20 cents, so it is \(\frac{2}{3}\)